six years ago,anna is twice as old as brita. six years hence, brita will be as old as anna is now while anna will be twice as old as she was six years ago. how old are they now?
Answer by ankor@dixie-net.com(22740)

six years ago, Anna is twice as old as Brita.
six years hence, Brita will be as old as Anna is now while Anna will be twice as old as she was six years ago.
how old are they now?
:
Let a = A's present age
Let b = B's present age
:
Write an equation for each statement:
:
"six years ago, Anna is twice as old as Brita."
a - 6 = 2(b - 6)
a - 6 = 2b - 12
a = 2b - 12 + 6
a = 2b - 6
:
"six years hence, Brita will be as old as Anna is now"
b + 6 = a
:
"while Anna will be twice as old as she was six years ago."
a + 6 = 2(a-6)
a + 6 = 2a - 12
6 + 12 = 2a - a
18 = a
A is 18 yrs now
:
b + 6 = 18
b = 18 - 6
B = 12 yrs now